a man invests one sum of money at 5% and a second sum, $2000 greater than the first sum at 6%. After one year his total income from investments is $1000.
Answer by checkley79(3341)

.05x+.06(x+2,000)=1,000
.05x+.06x+120=1,000
,11x=1,000-120
.1x=880
x=880/.11
x=$8,000 invested @ 6%.
8,000+2,000=$10,000 invested @ 6%
Proof:
.05*8,000+.06*10,000=1000
400+600=$1,000